An annuity is set up that will pay $ 1200 per year for eight years. What is the present value​ (PV) of this annuity given that the discount rate is 7​%?
Answer #1

for --PV(C2,C3,C4,C5,C6) B 1 0 .07|| ст А 1 2 RATE(i/y) | 3 NPER(n) 4 PMT 5 FV 6 TYPE 7 PV 8 Rate Per Period Total No.of peri

010 - 1 year Cash flows 2 1 $ 1,200.00 3 2 $ 1,200.00 3 $ 1,200.00 4 $ 1,200.00 5 $ 1,200.00 6 $ 1,200.00 7 $ 1,200.00 8 $ 1,

D10 fic =SUM(D2:09) 1 year 2 1 32 Cash flows 1200 1200 1200 1200 1200 1200 1200 1200 pv@7% =1/1.07 =C2/1.07 =C3/1.07 =C4/1.07